OpenStudy (anonymous):

Hi Everyone, I have to find k for the equation 1000=10xk^.5xk when I'm simplifying k^.5 x k does it become k^1.5 or 2k^1.5 thanks in advance :)

OpenStudy (dumbcow):

add the exponents .5 + 1 = 1.5 so the 1st one

OpenStudy (anonymous):

ha ha i over thought this question so bad. I thought he was writing 100=xk^(0.5xk) not 100=xk^(0.5)xk ha ha ha

OpenStudy (dumbcow):

yeah i can see that cj19, in future use parenthesis or * symbol for multiplication instead of x

OpenStudy (anonymous):

thanks everyone, I will set it out more clearly next time :)

OpenStudy (dumbcow):

thanks were you able to solve for k

OpenStudy (anonymous):

no worries its clear now because thinking it was and x^x situation is just silly. ;)

OpenStudy (anonymous):

I think so i got 21.54434, not too sure if that's correct though, in my workings I did \[\sqrt[1.5]{100}= k\]

OpenStudy (anonymous):

k^1.5